# BiLD: Bi-directional Logits Difference Loss for Large Language Model Distillation

### v2

@@ Abstract
In recent years, large language models (LLMs) have shown exceptional capabilities across various natural language processing (NLP) tasks. However, such impressive performance often comes with the trade-off of an increased parameter size, posing significant challenges for widespread deployment. Knowledge distillation (KD) provides a solution by transferring knowledge from a large teacher model to a smaller student model. In this paper, we explore the task-specific distillation of LLMs at the logit level. Our investigation reveals that the logits of fine-tuned LLMs exhibit a more extreme long-tail distribution than those from vision models, with hidden "noise" in the long tail affecting distillation performance. Furthermore, existing logits distillation methods often struggle to effectively utilize the internal ranking information from the logits. To address these, we propose the Bi-directional Logits Difference (BiLD) loss. The BiLD loss filters out the long-tail noise by utilizing only top-$k$ teacher and student logits, and leverages the internal logits ranking information by constructing logits differences. To evaluate BiLD loss, we conduct comprehensive experiments on 13 datasets using two types of LLMs. Our results show that the BiLD loss, with only the top-8 logits, outperforms supervised fine-tuning (SFT), vanilla KL loss, and five other distillation methods from both NLP and CV fields.
@@ 1 Introduction
Figure 1: An illustration of vanilla KL, top-$k$ KL and our BiLD loss. The vanilla KL loss directly calculates the KL divergence between teacher and student logits, whereas the top-$k$ KL loss uses clipped logits instead of the full logits. In contrast to these methods, our BiLD loss computes KL divergence based on reconstructed logits differences. The logits difference is derived by calculating the pairwise differences between logit values. We construct two groups of logits differences and compute the KL divergence within each group as a loss: the top-$k$ teacher logits and their corresponding student logits are used to calculate the teacher-led logits difference ($t$-LD) loss, while the top-$k$ student logits and their corresponding teacher logits are used to calculate the student-led logits difference (($s$-LD)) loss. The BiLD loss is the sum of these two losses.
The last few years have witnessed large language models (LLMs) risen to prominence, demonstrating remarkable proficiency in natural language understanding and generation. However, these capabilities come at the cost of an ever-increasing number of parameters. Due to constraints on computational resources, LLMs’ formidable size hinders their democratization and widespread deployment. Knowledge distillation (KD) , as a classic model compression method(Hinton et al., 2015), provides a solution for reducing model size while striving to maintain performance. KD transfers knowledge from a large teacher model to a smaller student model, thereby enhancing the student model’s performance and making it a lightweight alternative for deployment.
@@
For LLMs, research on logits distillation is still emerging, with methods such as reverse KL (Tu et al., 2020; Lee et al., 2023; Gu et al., 2023) and those based on optimal transport metrics (Cui et al., 2024). However, in practical applications, the former suffers from the "mode-seeking" problem (Chan et al., 2022; Li and Farnia, 2023), while the latter is computationally too complex for open-source large models with billions of parameters.
@@
Motivated by logits characteristics and the inadequate utilization of logits’ internal ranking, we design a novel loss, the Bi-directional Logits Difference (BiLD) loss, for task-specific LLM distillation. BiLD loss emphasizes reducing long-tail noise and explicitly utilizes the ranking information in logits. It computes KL divergence based on reconstructed logits differences, which are obtained by calculating the internal pairwise differences of values from top-$k$ teacher (student) logits and the corresponding student(teacher) logits. Our experiments show that BiLD loss, using only the top-8 logits, achieves state-of-the-art (SOTA) results across various NLP tasks. Our code is available at https://github.com/fpcsong/BiLD.
@@
  • We propose the Bi-directional Logits Difference (BiLD) loss for logits distillation of LLMs. BiLD filters out "noise" in logits’ long-tail distribution while leveraging logits ranking information to improve performance. Our method can serve as an alternative to the vanilla KL loss in existing LLM distillation methods.
@@ 2.1 Logits Distillation
Recent research has introduced several logit distillation methods suited for LLMs. Reverse KL (RKL) (Tu et al., 2020; Gu et al., 2023) has been used to mitigate the "mode-averaging" problem; however, it sometimes leads the student model towards "mode-seeking" behavior. DistiLLM (Ko et al., 2024) proposes mixing the logits distributions of the teacher and the student, but this introduces additional hyperparameters, increasing its complexity in practical applications. SinKD (Cui et al., 2024) replaces KL divergence with Sinkhorn Distance, but its computational demands pose challenges when applied to larger models.
@@ 3.2 The Characteristics of LLMs’ Logits
Compared to vision models, LLMs have an output space consisting of infinitely long sequences of tokens, making their logits more complex. We conduct a toy experiment to compare the logit characteristics of vision models and LLMs. We choose ResNet-101 (He et al., 2016) and Qwen-4B (Bai et al., 2023) for the toy case. We randomly select five images and five sets of instructions from our test data as inputs for the vision and language models (details about images and instructions are provided in Appendix C). We use kurtosis to measure the extremity of logits’ long-tail distribution and calculate the proportion of top-$k$ logit values. We report the experimental results in Table 1. The kurtosis of text logits is 2-3 orders of magnitude higher than that of image logits, suggesting that text logits are much "sharper" than image logits. Given that text logits are much longer than image logits, the proportion of top-$k$ logit values also indicates that text logit values are more concentrated than those of image logits.
@@ 3.3.1 Formal Definition
The BiLD loss consists of two components: the teacher-led logits difference ($t$-LD) loss and the student-led logits difference ($s$-LD) loss. Given the similarity between the two components, we explain the process using the calculation of $t$-LD loss. First, we select the top-$k$ teacher logits and sort them in descending order to build the teacher-led logits $\textbf{z}^{t}_{\rm led}$:
@@ 3.3.2 Explanation about the Utilization of Logits Ranking
The calculation of the logits difference (Equation 6) ensures that the student model learns the ranking information embedded in the teacher logits. We demonstrate this by taking the calculation of the $t$-LD loss as an example. Since $\textbf{z}^{t}_{\rm led}$ satisfies $z^{t}_{i_{1}}\geq z^{t}_{i_{2}}\geq\cdots\geq z^{t}_{i_{k}}$, it is guaranteed that every element in the teacher-led logits difference $\textbf{d}^{t}_{\rm led}$ is non-negative. For the corresponding student logits difference $\textbf{d}^{s}_{\rm cor}$, consider an element $d^{s}=z^{s}_{i_{m}}-z^{s}_{i_{n}}$. If $z^{s}_{i_{m}}<z^{s}_{i_{n}}$, then $d^{s}<0$. In this case, the order $z^{s}_{i_{m}}<z^{s}_{i_{n}}$ is inconsistent with the order in the teacher logits $z^{t}_{i_{m}}>z^{t}_{i_{n}}$. Therefore, the sign of the elements in the corresponding logits difference $\textbf{d}^{s}_{\rm cor}$ reflects whether the ranking of the teacher and student logits value pairs is consistent. When calculating $\mathcal{L}_{s-{\rm LD}}$, this acts as a constraint, promoting the student logits to align their ranking order with the teacher logits.
@@ 4.1 Datasets
We evaluate our BiLD loss on 13 NLP datasets: (1) 8 datasets from the SuperGLUE benchmark (Wang et al., 2019), including BoolQ (Clark et al., 2019), CB (De Marneffe et al., 2019), COPA (Roemmele et al., 2011), MultiRC (Khashabi et al., 2018), ReCoRD (Zhang et al., 2018), RTE (Giampiccolo et al., 2007), WiC (Pilehvar and Camacho-Collados, 2018) and WSC (Levesque et al., 2012); (2) 5 extra datasets used in previous works about model compression (Ma et al., 2023; Egiazarian et al., 2024), including: Arc-C, Arc-E (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020) and WinoGrande (Sakaguchi et al., 2021). We observe that these datasets vary significantly in size (the visualization of dataset sizes is presented in Appendix A). Using small datasets alone for SFT and distillation would result in severe overfitting. To prevent unreliable experimental results, we use these datasets collectively for SFT and distillation and evaluate each separately.
@@ 4.3 Implementation Details
We conduct experiments using the BLOOM and Qwen1.5 (abbreviated as Qwen) models, chosen for their availability in various sizes. Specifically, We employ BLOOM-7B and Qwen-4B as teacher models. For student models, we select BLOOM-3B and BLOOM-1B from the BLOOM series, and 1.8B and 0.5B versions from Qwen.
We perform three epochs of SFT on each teacher model and eight epochs of distillation for each student. Both SFT and distillation processes are conducted with a batch size of 64 and a micro batch size of 2, using the full dataset. We employ a cosine scheduler with an initial learning rate of $51e-$ 5 for SFT and $52e-$ 5 for distillation. The warm-up steps are set to 64. During SFT, we utilize the cross entropy loss. For the different distillation methods we tested, all parameters, except for temperature, are set to their default values. Due to the computational complexity of some distillation methods, we use the vanilla KL loss for the instruction part to expedite the distillation process, and apply different distillation losses to the output part. The temperature $T$ for all loss functions is set to 3. For the top-$k$ KL loss, we set $k$=1024, and for our proposed BiLD loss, we set $k$=8.
All our experiments are carried out on 8 NVIDIA A100 GPUs. To reduce memory usage, we employ DeepSpeed during both SFT and distillation processes, along with gradient checkpointing and BFLOAT16 mode (Kalamkar et al., 2019). We have not explored the minimum memory requirements. However, in practice, except for the DKD (Zhao et al., 2022), NKD (Yang et al., 2023b), and NormKD (Chi et al., 2023), experiments involving other methods can be conducted with half of the computational resources. During the evaluation, we employ vLLM (Kwon et al., 2023) for faster inference. The evaluation can be performed with a single NVIDIA A100 GPU. More implementation details can be found in our open-source repository.
@@ 5.1 Main Results
We report the experimental results on all 13 datasets in Table 2. Across all four sets of distillation, the BiLD loss achieves the highest average accuracy, outperforming SFT, vanilla KL, and the other five methods we tested. In the distillation from Qwen-4B to 0.5B, the BiLD loss showed a significant improvement in average accuracy, surpassing the vanilla KL loss by 3.52%. This improvement is also observed in the distillation from Qwen-4B to 1.8B and from BLOOM-7B to 1B, with improvements of 1.09% and 1.10% over the vanilla KL loss, respectively. A notable case is the distillation from BLOOM-7B to 1B, where the student using vanilla KL loss can easily match the teacher’s performance. In this scenario, our BiLD loss still maintained a consistent advantage, showing an average increase of 0.64% over the vanilla KL loss. In contrast, other methods achieve only marginal performance improvements or even experience declines. The robust performance of the BiLD loss across various distillation scenarios underscores its superiority and effectiveness.
@@ 5.2 Analysis of the Effectiveness of Clipping Logits
The experimental results in Table 2 indicate that in three sets of distillation, simply clipping the full logits to the top-$k$ logits improves the performance of the KL loss. This suggests that filtering out the noise in the logits’ long tail distribution can be a practical and straightforward approach to enhancing distillation performance. Our statistics show that the top-1024 logits cover over 99% of the probability in both Qwen-4B and BLOOM-7B teachers. For computational simplicity, we set $k$=1024 for the top-$k$ KL loss to verify that excluding the long-tail distribution of logits can improve distillation results.
@@ 5.3 Analysis of Performance at the Logit Level
where ${\rm topk(\cdot)}$ is a function to select tokens corresponding to the top-$k$ logit values. The metric ${\rm overlap}@{k}$ measures the average degree of overlap for the top-$k$ logits corresponding tokens at the same positions in $\textbf{Z}^{t}$ and $\textbf{Z}^{s}$. Specifically, ${\rm overlap}@{1}$ evaluates if the token corresponding to the highest logit values of both the teacher’s and the student’s outputs match at each position. ${\rm Overlap}@{1}$ can measure the efficacy of LLMs in greedy search mode, where LLMs generate text based on the token with the highest probability. For $k>1$, ${\rm overlap}@{k}$ calculates the ratio of overlapping tokens corresponding to the top-$k$ logits from both student and teacher at each position, reflecting how well the student imitates the important parts of teacher logits. From another perspective, ${\rm overlap}@{1}$ measures the performance of models in scenarios where there is only one correct answer, while ${\rm overlap}@{k}(k>1)$ reflects the degree of similarity between the student and teacher responses in open-ended scenarios.
@@ 6 Conclusion
Appendix A Details about Datasets
Figure 4: A visualization of the dataset sizes. There are significant size differences among the datasets, with the smallest datasets (CB, COPA, WSC) differing by three orders of magnitude from the largest dataset (ReCoRD).
Appendix B Calculation Efficiency of BiLD

### v3

@@ Abstract
In recent years, large language models (LLMs) have shown exceptional capabilities across various natural language processing (NLP) tasks. However, such impressive performance often comes at the cost of an increased parameter size, posing significant challenges for widespread deployment. To address this issue, knowledge distillation (KD) provides a solution by transferring knowledge from a large teacher model to a smaller student model. In this paper, we explore the task-specific distillation of LLMs at the logit level. Our investigation reveals that the logits of fine-tuned LLMs exhibit a more pronounced long-tail distribution compared to those of vision models, with hidden "noise" in the long tail affecting distillation performance. Furthermore, existing logits distillation methods often struggle to effectively utilize the internal ranking information from the logits. To address these, we propose the Bi-directional Logits Difference (BiLD) loss. The BiLD loss filters out the long-tail noise by utilizing top-$k$ teacher and student logits, and leverages the internal logits ranking information by constructing logits differences. To evaluate BiLD loss, we perform comprehensive experiments on 13 datasets with two types of LLMs. Our results show that the BiLD loss, with only the top-8 logits, outperforms supervised fine-tuning (SFT), vanilla KL loss, and five other distillation methods from both NLP and CV fields. Code is available at https://github.com/fpcsong/BiLD.
@@ 1 Introduction
Figure 1: An illustration of vanilla KL, top-$k$ KL and our BiLD loss. The vanilla KL loss directly calculates the KL divergence between teacher and student logits, whereas the top-$k$ KL loss uses clipped logits instead of the full logits. In contrast to these methods, our BiLD loss computes KL divergence based on reconstructed logits differences. The logits difference is derived by calculating the pairwise differences between logit values. We construct two groups of logits differences and compute the KL divergence within each group as a loss: the top-$k$ teacher logits and their corresponding student logits are used to calculate the teacher-led logits difference ($t$-LD) loss, while the top-$k$ student logits and their corresponding teacher logits are used to calculate the student-led logits difference ($s$-LD) loss. The BiLD loss is the sum of these two losses.
The last few years have witnessed large language models (LLMs) risen to prominence, demonstrating remarkable proficiency in natural language understanding and generation. However, these capabilities come at the cost of an ever-increasing number of parameters. Due to constraints on computational resources, LLMs’ formidable size limits their democratization and widespread deployment. Knowledge distillation (KD) , as a classic model compression method(Hinton et al., 2015), provides a solution for reducing model size while preserving performance. KD transfers knowledge from a large teacher model to a smaller student model, enhancing the latter’s performance and enabling its lightweight deployment.
@@
For LLMs, research on logits distillation is still emerging, with methods such as reverse KL (Tu et al., 2020; Lee et al., 2023; Gu et al., 2023) and those based on optimal transport metrics (Cui et al., 2024). However, in practical applications, the former suffers from the "mode-seeking" problem (Chan et al., 2022; Li and Farnia, 2023), while the latter is computationally too complex for open-source LLMs with billions of parameters.
@@
Motivated by the characteristics of logits and the underutilization of their internal ranking, we propose a novel loss function, called Bi-directional Logits Difference (BiLD) loss, for task-specific LLM distillation. BiLD loss emphasizes reducing long-tail noise and explicitly utilizes the ranking information in logits. It computes KL divergence based on reconstructed logits differences, which are obtained by calculating the internal pairwise differences of values from top-$k$ teacher (student) logits and the corresponding student (teacher) logits. Our experiments show that BiLD loss, using only the top-8 logits, achieves state-of-the-art (SOTA) results across various NLP tasks. Our code is available at https://github.com/fpcsong/BiLD.
@@
  • We propose the Bi-directional Logits Difference (BiLD) loss for logits distillation of LLMs. BiLD filters out "noise" in logits’ long-tail distribution while leveraging logits ranking information to enhance performance. Our method can serve as an alternative to the vanilla KL loss in existing LLM distillation methods.
@@ 2.1 Logits Distillation
Recent research has introduced several logit distillation methods specifically for LLMs. Reverse KL (RKL) (Tu et al., 2020; Gu et al., 2023) has been used to mitigate the "mode-averaging" problem; however, it sometimes leads the student model towards "mode-seeking" behavior. DistiLLM (Ko et al., 2024) proposes mixing the logits distributions of the teacher and the student, but this introduces additional hyperparameters, increasing its complexity in practical applications. SinKD (Cui et al., 2024) replaces KL divergence with Sinkhorn Distance, but its computational demands pose challenges when applied to larger models.
@@ 3.2 The Characteristics of LLMs’ Logits
Compared to vision models, LLMs have an output space consisting of infinitely long sequences of tokens, which makes their logits more complex. To compare the logit characteristics of vision models and LLMs, we conduct a toy experiment using ResNet-101 (He et al., 2016) and Qwen-4B (Bai et al., 2023). In this experiment, we randomly select five images and five sets of instructions from our test data as inputs for the vision and language models (details about images and instructions are provided in Appendix C). We use kurtosis to measure the extremity of logits’ long-tail distribution and calculate the proportion of top-$k$ logit values. The experimental results are reported in Table 1. The kurtosis of text logits is 2-3 orders of magnitude higher than that of image logits, suggesting that text logits are much "sharper" than image logits. Given that text logits are much longer than image logits, the proportion of top-$k$ logit values also indicates that text logit values are more peaked than those of image logits.
@@ 3.3.1 Formal Definition
The BiLD loss consists of two components: the teacher-led logits difference ($t$-LD) loss and the student-led logits difference ($s$-LD) loss. Due to the similarity between the two components, we explain the process by calculating the $t$-LD loss. First, we select the top-$k$ teacher logits and sort them in descending order to build the teacher-led logits $\textbf{z}^{t}_{\rm led}$:
@@ 3.3.2 Explanation about the Utilization of Logits Ranking
The calculation of the logits difference (Equation 6) ensures that the student model learns the ranking information embedded in the teacher logits. We demonstrate this through an example based on the $t$-LD loss calculation. Since $\textbf{z}^{t}_{\rm led}$ satisfies $z^{t}_{i_{1}}\geq z^{t}_{i_{2}}\geq\cdots\geq z^{t}_{i_{k}}$, it is guaranteed that every element in the teacher-led logits difference $\textbf{d}^{t}_{\rm led}$ is non-negative. For the corresponding student logits difference $\textbf{d}^{s}_{\rm cor}$, consider an element $d^{s}=z^{s}_{i_{m}}-z^{s}_{i_{n}}$. If $m>n$, then $d^{s}<0$. In this case, the order $z^{s}_{i_{m}}<z^{s}_{i_{n}}$ is inconsistent with the order in the teacher logits $z^{t}_{i_{m}}>z^{t}_{i_{n}}$. Therefore, the sign of the elements in the corresponding logits difference $\textbf{d}^{s}_{\rm cor}$ reflects whether the ranking of the teacher and student logits value pairs is consistent. When calculating $\mathcal{L}_{s-{\rm LD}}$, this acts as a constraint, promoting the student logits to align their ranking order with the teacher logits.
@@ 4.1 Datasets
We evaluate our BiLD loss on 13 NLP datasets: (1) 8 datasets from the SuperGLUE benchmark (Wang et al., 2019), including BoolQ (Clark et al., 2019), CB (De Marneffe et al., 2019), COPA (Roemmele et al., 2011), MultiRC (Khashabi et al., 2018), ReCoRD (Zhang et al., 2018), RTE (Giampiccolo et al., 2007), WiC (Pilehvar and Camacho-Collados, 2018) and WSC (Levesque et al., 2012); (2) 5 extra datasets used in previous works about model compression (Ma et al., 2023; Egiazarian et al., 2024), including: Arc-C, Arc-E (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020) and WinoGrande (Sakaguchi et al., 2021). We observe that these datasets vary significantly in size (the visualization of dataset sizes is presented in Appendix B). Using small datasets alone for SFT and distillation would result in severe overfitting. To prevent unreliable experimental results, we use these datasets collectively for SFT and distillation, then evaluate each separately.
@@ 4.3 Implementation Details
We conduct experiments using the BLOOM and Qwen1.5 (abbreviated as Qwen) models, which are available in various sizes. Specifically, we employ BLOOM-7B and Qwen-4B as teacher models. For student models, we select BLOOM-3B and BLOOM-1B from the BLOOM series, and 1.8B and 0.5B versions from Qwen.
We perform three epochs of SFT on each teacher model and eight epochs of distillation for each student. Both SFT and distillation are conducted with a batch size of 64 and a micro batch size of 2, using the full dataset. We employ a cosine scheduler with an initial learning rate of $51e-$ 5 for SFT and $52e-$ 5 for distillation. The warm-up steps are set to 64. During SFT, we utilize the cross entropy loss. For the different distillation methods we tested, all parameters, except for temperature, are set to their default values. Due to the computational complexity of some distillation methods, we use the vanilla KL loss for the instruction part to expedite the distillation process, and apply different distillation losses to the output part. The temperature $T$ for all loss functions is 3. For the top-$k$ KL loss, we set $k$=1024, and for our proposed BiLD loss, we set $k$=8.
All our experiments are carried out on 8 NVIDIA A100 GPUs. To reduce memory usage, we employ DeepSpeed (Rasley et al., 2020) during both SFT and distillation processes, along with gradient checkpointing and BFLOAT16 mode (Kalamkar et al., 2019). We have not explored the minimum memory requirements. However, in practice, experiments involving all methods except DKD (Zhao et al., 2022), NKD (Yang et al., 2023b), and NormKD (Chi et al., 2023) can be conducted with half of the computational resources. During the evaluation, we employ vLLM (Kwon et al., 2023) for faster inference. The evaluation can be performed with a single NVIDIA A100 GPU. More implementation details can be found in our repository.
@@ 5.1 Main Results
We report the experimental results on all 13 datasets in Table 2. Across all four sets of distillation, the BiLD loss achieves the highest average accuracy, outperforming SFT, vanilla KL, and the other five methods we tested. In the distillation from Qwen-4B to 0.5B, the BiLD loss showed a significant improvement in average accuracy, surpassing the vanilla KL loss by 3.52%. This improvement is also observed in the distillation from Qwen-4B to 1.8B and from BLOOM-7B to 1B, with improvements of 1.09% and 1.10% over the vanilla KL loss respectively. A notable case is the distillation from BLOOM-7B to 1B, where the student using vanilla KL loss can easily match the teacher’s performance. In this scenario, our BiLD loss still maintained a consistent advantage, with an average increase of 0.64% over the vanilla KL loss. In contrast, other methods show only marginal improvements or even declines in performance. The robust performance of the BiLD loss across various distillation scenarios underscores its superiority and effectiveness.
@@ 5.2 Analysis of the Effectiveness of Clipping Logits
The experimental results in Table 2 indicate that, in three distillation scenarios, simply clipping the full logits to the top-$k$ logits improves the performance of the KL loss. This suggests that filtering out the noise in the logits’ long tail distribution can be a practical and straightforward approach to enhancing distillation performance. Our statistics show that the top-1024 logits cover over 99% of the probability in both Qwen-4B and BLOOM-7B teachers. For computational simplicity, we set $k$=1024 for the top-$k$ KL loss to verify that excluding the long-tail distribution of logits can improve distillation results.
@@ 5.3 Analysis of Performance at the Logit Level
where ${\rm topk(\cdot)}$ is a function to select tokens corresponding to the top-$k$ logit values. The metric ${\rm overlap}@{k}$ measures the average degree of overlap for the top-$k$ logits corresponding tokens at the same positions in $\textbf{Z}^{t}$ and $\textbf{Z}^{s}$. Specifically, ${\rm overlap}@{1}$ evaluates whether the tokens with the highest logit values from the teacher and student outputs match at each position. This metric can measure the efficacy of LLMs in greedy search mode, where LLMs generate text based on the token with the highest probability. For $k>1$, ${\rm overlap}@{k}$ calculates the ratio of overlapping tokens corresponding to the top-$k$ logits from both student and teacher at each position, reflecting how well the student imitates the important parts of teacher logits. From another perspective, ${\rm overlap}@{1}$ measures the performance of models in scenarios where there is only one correct answer, while ${\rm overlap}@{k}(k>1)$ reflects the degree of similarity between the student and teacher responses in open-ended scenarios.
@@ 6 Conclusion
Appendix A Calculation Efficiency of BiLD
@@
Appendix B Details about Dataset Sizes
Details about the dataset sizes are shown in Figure 5. There are significant size differences among the datasets, with the smallest datasets (CB, COPA, WSC) differing by three orders of magnitude from the largest dataset (ReCoRD).
Figure 5: A visualization of the dataset sizes.